# A Minimal Model for Emergent Collective Behaviors in Autonomous Robotic Multi-Agent Systems

### v1

@@ 1 Introduction
Collective behavior in biological systems, from bacteria to vertebrates, shares striking similarities in the emergence of global patterns and information transfer mechanisms, as seen in bird flocks [1], swarms of flying insects [2], and fish swarms or polarized schools [3]. These fascinating natural phenomena emerge from entities collectively, following relatively simple, decentralized mechanisms of interaction with their immediate neighbors within a limited range. Extensive scientific efforts have been devoted to uncovering the underlying mechanisms or rules driving these dynamics of collective motion. Reynolds [4] proposed three types of local interaction rules, namely collision avoidance or separation, velocity matching or alignment, and flock centering or cohesion, to imitate the aggregate motion of a flock of birds. The rules defined the behavior of a computational model, known as boids, whose collective motion exhibited realistic and complex emergent phenomena such as flocking, swarming, and evading predators.
@@
Following the Reynolds behavioral model, efforts have been made to mathematically model abstract collective biological behaviors. The notable models dedicated to defining local interaction rules from which complex behaviors emerge are the Vicsek model [5], zone-based interaction model [6], Olfati-Saber model [7], and Cucker-Smale model [8]. Many variations of the Vicsek model for particle systems have been proposed to mimic the properties of natural creatures [11, 10, 9, 12]. The Vicsek model describes self-propelled particles moving at a constant velocity, aligning their direction with the average orientation of nearby particles, and transitioning from disorder to order by tuning angular noise and density terms. The zone-based interaction model organizes local interactions into three zones: repulsion, orientation, and attraction, enabling the replication of collective behaviors through zone adjustments. Some variants of this model have been proposed to control swarm motion [13, 14]. Olfati-Saber introduced three types of agents: the $\alpha$-agents (cooperative agents), $\beta$-agents (obstacles), and $\gamma$-agents (leaders). Pairwise potential-based control laws are designed for $\alpha$-agent interactions to form $\alpha$-lattices, typically hexagonal, and to manage interactions with obstacles and leaders. The $\alpha$-lattice flock formation has gained significant attention in control applications [15, 16, 17]. The Cucker-Smale model focused on the alignment of velocity influenced by relative positions, where the alignment strength is dependent on distance. This model has been the subject of extensive study, with numerous extensions and variants proposed [18, 19, 20].
@@
Nevertheless, the interaction rules in the aforementioned models and their variations often fail to replicate the properties that make Reynolds’ boids excel in producing complex collective behaviors. For instance, the Vicsek model, zone-based interaction model, and Cucker-Smale model lack critical features such as flock centering, inter-agent collision avoidance, or obstacle avoidance. Meanwhile, the Olfati-Saber model tends to produce rigid and overly regular patterns, missing the flexibility inherent in the boids model.
@@
In particular, the ability to avoid collisions with neighboring entities or to maintain personal space, that is to say, to separate oneself from others, is a fundamental characteristic of biological swarms, such as bird flocks, as well as of robot swarms. However, the aforementioned models of swarms and flocks in their original form are unable to directly address the issue of avoiding collisions between neighbors while simultaneously exhibiting the spatial flexibility observed in nature. While numerous extensions to classical models employ repulsive potentials to implement collision avoidance, these approaches often come at the expense of increased model complexity, compromising their minimalistic nature. This highlights a significant research gap with regard to the development of innovative models that are capable of effectively incorporating this feature while preserving minimalism and computational efficiency.
@@
The remainder of this paper is organized as follows. Section 2 provides an overview of emergent collective behaviors and models. Sections 3 and 4 describe the model and its resource-aware cognitive extension, respectively. Section 5 examines the asymptotic stability of the proposed framework. Section 6 provides simulation results along with model analyses. Finally, Section 7 offers concluding remarks and outlines potential directions for future work.
@@ 2 Emergent Collective Behaviors and Models
Flocking, vortexing, and swarming are widespread emergent collective behaviors observed in biological groups. These behaviors emerge from individuals following simple rules through local interactions. Table 1 provides a comparative summary of their characteristics.
Table 1: Differences between flocking, vortexing, and swarming behaviors.
Behavioral aspect Flocking Vortexing Swarming
Coordination High Moderate Less
@@
To model these behaviors, animals are typically treated as simple entities governed by microscopic rules. These rules, which are in principle inspired by the sociological behavior of animals, are based on their position and velocity within a given neighborhood, as well as local density. The social tendencies for grouping (attraction/aggregation), the minimal space required for movement and comfort (repulsion/collision avoidance), and the ability to synchronize or align with others (orientation/alignment) form the basis of these rules [31]. In summary, these behaviors can be effectively characterized and distinguished by assessing aggregation and alignment in collective motion.
Flocking is characterized by both aggregation and alignment of the motion direction. This behavior enables coordinated group movements, as seen in flocks of migrating birds [32]. The flocking behavior emerges from the alignment of individual velocities and positions, facilitating synchronized group travel and efficient navigation over long distances.
Vortexing behavior, by which individuals turn around a common center, is commonly known by a variety of names, including ant mill, caterpillar circle, bat doughnut, amphibian vortex, duck swirl, and fish torus [33]. The potential benefit of collective vortex behaviors to individuals, for example, is protection against predators. This behavior involves elements of both local structure (e.g., agents moving in circular or spiral patterns) and global disorder (e.g., lack of overall alignment or uniform motion). Vortexing is structured in the sense that it includes repeating patterns (e.g., the vortex structure). It is disordered due to the absence of global alignment. Vortexing represents a moderate level of coordination, where agents maintain localized interactions to sustain circular patterns without achieving global alignment.
Swarming is characterized by aggregation and lack of overall alignment, where agents exhibit aggregate movement without coordinating their motion directions. This behavior is often driven by group objectives, such as search, coverage, foraging, or predator evasion [34]. Swarming typically results in a disordered yet cohesive structure, allowing groups to respond flexibly to environmental challenges while maintaining proximity among members.
In 1987, Reynolds [4] introduced three rules that govern the flocking behavior of birds. These rules are as follows: $i)$ collision avoidance: avoid collisions with nearby flockmates; ii)ii) velocity matching: attempt to match velocity with nearby flockmates; $iii)$ flock centering: attempt to stay close to nearby flockmates. Olfati-Saber [7] sought to express these behavioral rules mathematically. However, the resultant $\alpha$-lattice formation lacked a key property of the boids flocking behavior, a flexible and adaptable spatial configuration. The boids behavioral rules do not dictate a rigid lattice-type spatial formation. Vicsek [5] and Cucker-Smale [8] models have been widely acclaimed for showing cohesive flocking and consensus formation. Nevertheless, both models originally lacked flock centering, inter-agent separation, or obstacle avoidance. A comparative overview of the behavioral attributes of these models is presented in Table 2.
Table 2: Emergent behaviors of classical and proposed models.
Model Flocking behavior Swarming behavior Collision avoidance
@@
    Reynold’s model is heuristic and rule-based.
@@
    Olfati-Saber’s model ensures collision avoidance via explicit repulsive potentials—yielding strong separation.
@@
    The proposed model relies on implicit collision avoidance through an offset-vector-based repulsion mechanism—yielding soft yet effective separation.
The Vicsek model is based on the concept that a bird maintains a constant velocity magnitude and adjusts its heading (or angular velocity) to match the average heading of its neighboring birds. Specifically, each bird updates its heading by averaging the headings of neighboring birds within a predetermined distance, with control exerted over the angular velocity. Random noise is introduced to prevent perfect alignment, resulting in dynamic, non-deterministic interactions that can exhibit swarming behavior. The Cucker-Smale model relaxes the assumption of constant velocity magnitude in Vicsek’s model, allowing for better realism.
The Cucker-Smale model describes the flocking behavior through velocity alignment dynamics. Specifically, each agent adjusts its velocity by incorporating a weighted average of the velocity differences between itself and other agents.
Consider $n$ interacting agents ($n\geq 2$) navigating an $m$-dimensional space, where $m=2$ or $3$. For each agent indexed by $i=1,\ldots,n$, the position and velocity vectors are denoted by $\mathbf{p}_{i}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}\in\mathbb{R}^{m}$, respectively.
The Cucker-Smale model is governed by
@@
with constants $K,\sigma>0$, $\gamma\geq 0$, and $\lVert\cdot\rVert$ represents the Euclidean norm. The weight function depends on the distance between birds, influencing the strength of their velocity alignment.
@@
The proposed model may be viewed as a generalization of the Cucker-Smale flocking model, incorporating several notable distinctions. First, while the Cucker-Smale model primarily emphasizes alignment-based flocking behavior, our model introduces a spatial offset tuning parameter that regulates inter-agent distances, thereby implicitly promoting collision avoidance. Second, in contrast to the Cucker-Smale model’s distance-based weight function for velocity alignment, our model employs a relative speed-based influence function, governed by the kinetic offset parameter. When the kinetic offset is set to zero or a small value, agents converge to similar velocities, resulting in flocking behavior. Conversely, larger kinetic offset values allow agents to maintain nonzero relative velocities, facilitating swarming behavior.
@@ 3 The Proposed Collective Behavior Model
$\psi(\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert)=1-(\frac{\delta_{i}|\mathcal{N}_{i}|}{\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert})^{\alpha},$
$\phi(\lVert\mathbf{v}_{j}-\mathbf{v}_{i}\rVert)=1-(\frac{\eta_{i}}{|\mathcal{N}_{i}|\lVert\mathbf{v}_{j}-\mathbf{v}_{i}\rVert})^{\beta},$
where $\delta_{i},\eta_{i}\geq 0$ represent the spatial offset and kinetic offset, respectively, $|\mathcal{N}_{i}|$ denotes the cardinality of the set $\mathcal{N}_{i}$, and $\alpha,\beta>0$.
The proposed model regulates agent interactions via two core mechanisms: an aggregation potential and an alignment potential. Within this framework, the spatial offset $\delta_{i}$ and kinetic offset $\eta_{i}$ govern the degrees to which an agent engages in aggregation and alignment with its neighbors, respectively. The parameters $\alpha$ and $\beta$ are introduced to maintain dimensional consistency, ensuring the model’s physical plausibility. Together, these design elements modulate each agent’s sensitivity to local interactions, giving rise to the emergent collective dynamics.
The aggregation potential induces the cohesion-separation interaction and dictates group spacing. The potential function minimum defines the equilibrium inter-agent distance. Here, $\delta_{i}$ scales this aggregation potential, which combines short-range repulsion with long-range attraction. A smaller $\delta_{i}$ enhances attractive forces (while preserving repulsive responses at close range), resulting in a denser formation. Conversely, a larger $\delta_{i}$ reduces attraction, promoting looser spacing. Therefore, $\delta_{i}$ effectively sets the equilibrium spacing of the collective, enforcing implicit collision avoidance.
In parallel, $\eta_{i}$ modulates the velocity-alignment coupling, determining how strongly an agent adjusts its velocity to match that of its neighbors. A smaller $\eta_{i}$ increases responsiveness to velocity differences, fostering rapid convergence toward a unified heading and supporting coherent flocking behavior. A larger $\eta_{i}$, by contrast, reduces this coupling, allowing for persistent velocity discrepancies and enabling the emergence of more complex collective behaviors, such as swarming or, in smaller populations, vortexing.
Thus, $\delta_{i}$ and $\eta_{i}$ play a central role in modulating the emergent dynamics of the collective system. These parameters serve as tunable levers for enabling transitions between distinct collective behaviors based on environmental demands or internal agent states. Specifically, $\delta_{i}$ scales the gradient of the inter-agent aggregation potential that pulls agents together when they are far apart and pushes them apart when they get too close, thereby regulating spatial cohesion and enforcing implicit collision avoidance by establishing the equilibrium inter-agent spacing that functions as a safe distance. Similarly, $\eta_{i}$ modulates the gradient of the alignment potential, enhancing convergence to velocity consensus when small, or sustaining velocity diversity when large. Overall, $\delta_{i}$ governs the spatial aggregation (and separation or implicit collision-avoidance) behavior, $\eta_{i}$ governs the temporal alignment (or misalignment) behavior, and $\alpha$, $\beta$ adjust interaction sensitivity. Tuning these parameters enables transitions between coherent flocking and loosely coordinated swarming, and should be guided by the target application and desired collective behavior. The roles of these parameters in shaping the emergent collective behaviors are summarized in Table 3.
Table 3: Roles of model parameters in shaping collective dynamics
@@
$\delta\downarrow$ Stronger, denser aggregation—characteristic of flocking
$\delta\uparrow$ Weaker spatial cohesion—characteristic of swarming
$\eta\downarrow$ Velocity alignment—characteristic of flocking
$\eta\uparrow$ Velocity misalignment—characteristic of swarming
$\alpha,\beta\uparrow$ Steeper responses to spatial or velocity differences
$\alpha,\beta\downarrow$ Smoother responses to spatial or velocity differences
The proposed model is inherently scalable, as each agent interacts only with neighbors within a fixed range $r_{i}$. In large, spatially distributed systems, the neighbor set size $|\mathcal{N}_{i}|$ typically remains bounded, ensuring that the per-agent computational complexity is independent of the total number of agents $n$. This locality makes the model well-suited for large-scale robotic swarms.
@@
The resulting displacement offset vector quantifies the repulsion or attraction tendency of an agent relative to its neighbors, thereby shaping the spatial structure and cohesion of the group. Similarly, the velocity difference offset vector governs the agent’s alignment behavior, with large magnitudes promoting directional heterogeneity—resulting in swarming or vortex-like collective dynamics. In contrast, small or vanishing velocity offsets encourage synchronization of velocities, which facilitates the coherent and aligned group motion characteristic of flocking behavior.
Implicit collision avoidance is intrinsically embedded in the model via the displacement offset vector-based repulsion mechanism. When a neighboring agent $j$ approaches agent $i$, the resulting repulsive interaction drives the distance between agents $\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert$ toward an equilibrium value determined by spatial offset. At this equilibrium, the repulsion is attenuated. Mathematically, the interaction weight $\psi(\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert)$ becomes increasingly negative as $\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert\to 0$, thereby inducing a stronger repulsive force that pushes agent $i$ away from agent $j$. As a result, the model inherently enforces a minimum separation distance, ensuring collision avoidance without the need for additional constraints or supervisory mechanisms.
In the local interaction model defined in (3), each agent modulates its behavior based on both the topological structure of its local neighborhood $\mathcal{N}_{i}$ and the metric properties of relative positions and velocities. This dual dependence enables a robust framework for decentralized and distributed control in robotic multi-agent systems. In particular, the inclusion of the neighborhood cardinality $|\mathcal{N}_{i}|$ in the modulation of the spatial and kinetic offsets $\delta_{i}$ and $\eta_{i}$ ensures that the influence of neighbors is adaptively scaled relative to the density of the local agent. This dynamic scaling mechanism promotes flexible interaction responses under varying crowd conditions and prevents individual agents from being overwhelmed by excessive neighbor influence. Consequently, the model fosters stable and coherent collective dynamics, enhancing its suitability for large-scale and density-varying environments.
The model’s ability to modulate interaction strengths based on real-time local density renders it particularly advantageous for real-world applications in swarm robotics and multi-agent simulations, where agent populations and spatial distributions are often dynamic. Such scalability ensures responsiveness and stability in complex and unstructured operational settings.
In summary, the minimalistic interaction model (3) is capable of generating a rich repertoire of collective behaviors. It establishes a balance between repulsion, which prevents overcrowding, and attraction, which promotes group cohesion. By appropriately tuning the offset parameters $\delta_{i}$ and $\eta_{i}$, one can elicit a diverse range of emergent behaviors—such as flocking and swarming across varying group sizes and vortexing in small-scale groups.

### v2

@@ 1 Introduction
Collective behavior in biological systems, from bacteria to vertebrates, exhibits striking similarities in emergent global patterns and information transfer, as observed in bird flocks [1], insect swarms [2], and fish schools [3]. These phenomena arise from simple, decentralized interactions among entities within a limited range. Reynolds [4] introduced three local interaction rules: separation, alignment, and cohesion, in his boids model, which simulates realistic collective motion like flocking and predator evasion.
@@
Subsequent efforts have mathematically modeled collective behaviors. The notable models include the Vicsek model [5], zone-based interaction model [6], Olfati-Saber model [7], and Cucker-Smale model [8]. Many variations of the Vicsek model for particle systems have been proposed to mimic the properties of natural creatures [11, 10, 9, 12]. The Vicsek model describes self-propelled particles moving at a constant velocity, aligning their direction with the average orientation of nearby particles, and transitioning from disorder to order by tuning angular noise and density terms. The zone-based interaction model organizes local interactions into three zones: repulsion, orientation, and attraction, enabling the replication of collective behaviors through zone adjustments. Some variants of this model have been proposed to control swarm motion [13, 14]. Olfati-Saber introduced three types of agents: the $\alpha$-agents (cooperative agents), $\beta$-agents (obstacles), and $\gamma$-agents (leaders). Pairwise potential-based control laws are designed for $\alpha$-agent interactions to form $\alpha$-lattices, typically hexagonal, and to manage interactions with obstacles and leaders. The $\alpha$-lattice flock formation has gained significant attention in control applications [15, 16, 17]. The Cucker-Smale model focused on the alignment of velocity influenced by relative positions, where the alignment strength is dependent on distance. This model has been the subject of extensive study, with numerous extensions and variants proposed [18, 19, 20].
@@
Nevertheless, the interaction rules in these models and their variations often fail to replicate the properties that make Reynolds’ boids excel in producing complex collective behaviors. For instance, the Vicsek model, zone-based interaction model, and Cucker-Smale model lack critical features such as flock centering, inter-agent collision avoidance, or obstacle avoidance. Meanwhile, the Olfati-Saber model tends to produce rigid and overly regular patterns, missing the flexibility inherent in the boids model.
The ability to avoid collisions and maintain personal space is fundamental in both biological and robotic swarms. Classical models do not inherently address this aspect of collective motion without incorporating complex repulsive potentials, thereby compromising their minimalism. This highlights a research gap for innovative models that effectively incorporate collision avoidance while preserving simplicity and computational efficiency.
@@
The remainder of this paper is organized as follows. Section 2 provides an overview of emergent collective behaviors and models. Sections 34, and 5 describe the model and its extensions incorporating target-directed and obstacle avoidance potentials to enable navigation toward global objectives in cluttered environments, and resource-aware cognitive extension, respectively. Section 6 examines the asymptotic stability of the proposed framework. Section 7 provides simulation results along with model analyses. Finally, Section 8 offers concluding remarks and outlines potential directions for future work.
@@ 2 Emergent Collective Behaviors and Models
Flocking, vortexing, and swarming are emergent behaviors in biological groups, driven by simple local interaction rules. Table 1 summarizes their characteristics.
Table 1: Characteristics of flocking, vortexing, and swarming.
Aspect Flocking Vortexing Swarming
Coordination High Moderate Low
@@
These behaviors arise from rules based on position, velocity, and local density, balancing attraction, repulsion, and alignment [31]. Flocking involves strong alignment and aggregation for coordinated movement, as in migrating birds [32]. Vortexing features agents circling a center, balancing local structure and global disorder for protection [33]. Swarming emphasizes aggregation without alignment, enabling flexible responses for foraging or evasion [34].
Reynolds’ boids model [4] uses collision avoidance, velocity matching, and flock centering. Olfati-Saber’s model [7] formalizes these but yields rigid lattices, unlike boids’ flexibility. Vicsek [5] and Cucker-Smale [8] models enable flocking but lack collision avoidance. Our model achieves flexible, collision-free swarming and flocking, overcoming Vicsek and Cucker-Smale’s limitations and Olfati-Saber’s rigidity, enhancing swarm robotics applicability. Table 2 compares these models.
Table 2: Emergent behaviors of models.
Model Flocking Swarming Collision Avoidance
@@
    Heuristic, rule-based model.
@@
    Uses repulsive potentials for strong separation.
@@
    Employs offset-vector-based repulsion for soft, effective separation.
Consider $n$ interacting agents ($n\geq 2$) navigating an $m$-dimensional space, where $m=2$ or $3$. For each agent indexed by $i=1,\ldots,n$, the position and velocity vectors are denoted by $\mathbf{p}_{i}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}\in\mathbb{R}^{m}$, respectively. The Cucker-Smale model describes the flocking behavior through velocity alignment dynamics. Specifically, each agent adjusts its velocity by incorporating a weighted average of the velocity differences between itself and other agents. The Cucker-Smale model is governed by
@@
with constants $K,\sigma>0$, $\gamma\geq 0$, and $\lVert\cdot\rVert$ represents the Euclidean norm. The weight function depends on the distance between agents, influencing the strength of their velocity alignment.
Extensions of the Cucker-Smale model address its lack of collision avoidance. For instance, interparticle bonding forces were introduced to ensure collision-free flocking [35], while sharp conditions for avoiding collisions in singular Cucker-Smale interactions were established for general initial configurations [36]. Notably, collision avoidance without additional forces was proven for some singular short-range-communicated Cucker-Smale models [37].
@@
The proposed model may be viewed as a generalization of the Cucker-Smale flocking model, incorporating several notable distinctions. First, while the Cucker-Smale model primarily emphasizes alignment-based flocking behavior, our model introduces a spatial offset tuning parameter that regulates inter-agent distances, thereby implicitly promoting collision avoidance. Second, in contrast to the Cucker-Smale model’s distance-based weight function for velocity alignment, our model employs a relative speed-based influence function, governed by the kinetic offset parameter. A low kinetic offset yields flocking via velocity consensus, while a high offset enables swarming through sustained velocity differences.
Remark 2.
The kinetic offset functions similarly to noise in the Vicsek model: smaller values promote alignment, like low noise, while larger values allow disordered motion, akin to high noise levels.
@@ 3 The Proposed Collective Behavior Model
$\psi(\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert)=1-(\frac{\delta_{i}|\mathcal{N}_{i}|}{\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert})^{\alpha_{i}},$
$\phi(\lVert\mathbf{v}_{j}-\mathbf{v}_{i}\rVert)=1-(\frac{\eta_{i}}{|\mathcal{N}_{i}|\lVert\mathbf{v}_{j}-\mathbf{v}_{i}\rVert})^{\beta_{i}},$
where $\delta_{i},\eta_{i}\geq 0$ are the spatial and kinetic offsets, respectively, $|\mathcal{N}_{i}|$ denotes the cardinality of the set $\mathcal{N}_{i}$, and $\alpha_{i},\beta_{i}>0$ control the steepness of the response.
The model balances aggregation (cohesion-separation) and alignment potentials. The spatial offset $\delta_{i}$ regulates inter-agent spacing, promoting denser formations when small and looser spacing when large, implicitly ensuring collision avoidance. The kinetic offset $\eta_{i}$ controls velocity alignment: smaller values enhance consensus for flocking, while larger values allow velocity diversity for swarming or vortexing. Offsets $\delta_{i}$ and $\eta_{i}$ tune transitions between flocking and swarming, with $\alpha_{i}$ and $\beta_{i}$ adjusting interaction sensitivity. Weight function profiles for $\psi$ and $\phi$ are shown in Figs. 1(1(a)) and 1(1(b)). Table 3 summarizes their effects.
Figure 1: (a) Aggregation weight $\psi$ for $\delta_{i}=0.25,0.5,0.75,1$, $\alpha=1$, $|\mathcal{N}_{i}|=5$. Positive $\psi$ induces cohesion, negative $\psi$ induces separation, zero at $\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert=\delta_{i}|\mathcal{N}_{i}|$. (b) Alignment weight $\phi$ for $\eta_{i}=0,1,2,5$, $\beta=2$, $|\mathcal{N}_{i}|=5$. Positive $\phi$ promotes alignment, negative $\phi$ reduces alignment, zero at $\lVert\mathbf{v}_{j}-\mathbf{v}_{i}\rVert=\frac{\eta_{i}}{|\mathcal{N}_{i}|}$.
Table 3: Effects of model parameters on collective dynamics.
@@
$\delta\downarrow$ Denser aggregation (flocking)
$\delta\uparrow$ Looser spacing (swarming)
$\eta\downarrow$ Stronger velocity alignment (flocking)
$\eta\uparrow$ Velocity misalignment (swarming)
$\alpha,\beta\uparrow$ Steeper interaction responses
$\alpha,\beta\downarrow$ Smoother interaction responses
@@
The displacement offset vector drives repulsion or attraction, enforcing equilibrium spacing at $\delta_{i}|\mathcal{N}_{i}|$, while the velocity offset vector governs alignment, enabling flocking or swarming based on $\eta_{i}$. Implicit collision avoidance is embedded via the repulsion mechanism, where the interaction weight $\psi(\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert)$ becomes strongly negative as $\lVert\mathbf{p}_{j}-\mathbf{p}_{i}\rVert\to 0$, pushing agents apart to maintain a minimum separation distance without additional constraints.
The model is scalable, as each agent interacts only with neighbors within $r_{i}$, keeping computational complexity independent of $n$. Agents adjust behavior based on local neighborhood $\mathcal{N}_{i}$ and relative positions/velocities. Including $|\mathcal{N}_{i}|$ in spatial ($\delta_{i}$) and kinetic ($\eta_{i}$) offsets adaptively scales neighbor influence by density. The separation equilibrium distance $|\mathcal{N}_{i}|\delta_{i}$ increases in crowded settings for more space and decreases in sparse ones for cohesion. This ensures stable, flexible dynamics in large-scale, varying-density environments.
4 Target-Directed Collective Behavior in Cluttered Environments
To account for migration toward a preferred target location and obstacle avoidance in cluttered environments, we extend the collective behavior model introduced in Section 3. These extensions are inspired by natural collectives, where agents orient their motion toward a common goal while avoiding collisions with environmental obstacles. For robotic multi-agent systems, the interaction rules incorporate: ($i$) migration toward a target location, which steers agents toward a shared objective; ($ii$) obstacle avoidance, which prevents collisions with environmental obstacles; and (iiiiii) the existing aggregation and alignment interactions, which regulate inter-agent cohesion, separation, and velocity alignment.
The target location, denoted by $\mathbf{p}_{t}\in\mathbb{R}^{m}$, is known to all agents and establishes a common direction for the collective. The environment contains a set of $O$ obstacles, denoted by $\mathcal{O}=\{1,\dots,O\}$, with each obstacle $o\in\mathcal{O}$ located at position $\mathbf{p}_{o}\in\mathbb{R}^{m}$. The set of obstacles detected by agent $i$ is defined as $\mathcal{O}_{i}=\{o\in\mathcal{O}\mid\lVert\mathbf{p}_{o}-\mathbf{p}_{i}\rVert\leq c_{i}\}$, where $c_{i}>0$ is the obstacle detection/interaction radius of agent $i$.
The extended dynamics of agent $i$ in a cluttered environment with a target are governed by the following interaction model, which augments the original dynamics in (3):
$\frac{d\mathbf{p}_{i}}{dt}$ $=\mathbf{v}_{i},$
$\frac{d\mathbf{v}_{i}}{dt}$ $=\text{aggregation potential}+\text{alignment potential}$
$\quad+\underbrace{\kappa_{i}(\mathbf{p}_{t}-\mathbf{p}_{i})}_{\text{target potential}}+\underbrace{\sum_{o\in\mathcal{O}_{i}}\rho(\lVert\mathbf{p}_{o}-\mathbf{p}_{i}\rVert)(\mathbf{p}_{o}-\mathbf{p}_{i})}_{\text{obstacle avoidance potential}},$ (3)
where $\kappa_{i}>0$ is a constant weight controlling the strength of attraction to the target, and the obstacle avoidance interaction weight is defined as
$\rho(\lVert\mathbf{p}_{o}-\mathbf{p}_{i}\rVert)=\min(0,\,1-(\frac{c_{i}}{\lVert\mathbf{p}_{o}-\mathbf{p}_{i}\rVert})^{\sigma_{i}}),$
with $\sigma_{i}>0$ controlling the steepness of the repulsion response.
The obstacle avoidance interaction weight $\rho$ is designed to induce repulsion from obstacles, analogous to the repulsion component of the aggregation interaction weight $\psi$. The function $\rho$ decreases sharply as the distance $\lVert\mathbf{p}_{o}-\mathbf{p}_{i}\rVert$ approaches zero, generating a strong repulsive force to steer agent $i$ away from obstacle $o$. This behavior ensures effective collision avoidance. The profile of $\rho$ is illustrated in Fig. 2.
Figure 2: Obstacle avoidance weight $\rho$ for $\sigma=2,1.5,1,0.5,0.25$ at detection/interaction radius $c_{i}=10$.